# Euler characteristic of crepant resolutions of specific modular quotient singularities

### v1

@@ 4 Non-abelian case of the main theorem
Let $G$ be a non-abelian group with a semidirect product structure $H\rtimes C_{p}$. Then there does not exist any Galois extension $L/K$ such that $\mathrm{Gal}(L/K)\cong G$. In particular, $f_{G}=0$.
@@
If there exists such a Galois extension $L/K$, then $L^{C(G)}/K$ is a Galois extension with Galois group isomorphic to $G/C(G)\cong(H/C(G))\rtimes C_{p}$. Therefore, we can assume $C(G)=\{e\}$ without loss of generality.
@@
By classical results (one can refer to [serre2013local], IV, Corollary 4), $G_{0}$, as a normal subgroup of $G$, is a semidirect product of a normal $p$-subgroup and a non-modular cyclic group. If $G_{0}$ contains an element of order $p$, then $G_{0}$ has to be the whole $G$ since $G_{0}\triangleleft G$, which contradicts to the semidirect product structure of $G_{0}$. Hence $C_{p}$ is not contained in $G_{0}$. Since $G_{0}=\{g\in G\mid g\text{ acts trivially on }\mathrm{res}(L)\}$, $C_{p}$ acts effectively on $\mathrm{res}(L)$.
Hereby, for $L^{H}/K$, as an Artin-Schreier extension which is either unramified or totally ramified, its residue degree $f_{L^{H}/K}>1$. Consequently, $L^{H}/K$ is unramified. Therefore, if we write the local field $L\cong\mathrm{res}(L)((t_{L}))$, there exists an element $g\in G$ of order $p$ acting trivially on $t_{L}$. We assume that $C_{p}$ is generated by such $g$ without loss of generality.
Take an element $\sigma$ of prime order $l$ ($l\neq p$) from $H$. Then $L/L^{<\sigma>}$ is studied by Kummer theory. If $L/L^{<\sigma>}$ is unramified, then we can similarly assume that $\sigma$ acts trivially on $t_{L}$, and thus $\sigma$ has to commute with $C_{p}$, since Galois groups of extensions of finite fields are abelian. If $L/L^{<\sigma>}$ is totally ramified, then $\sigma$ acts trivially on $\mathrm{res}(L)$, and hence also commutes with $C_{p}$.
@@
Lemma 4.1 only holds in equal characteristic. In mixed characteristic, for example, the splitting field of $x^{13}+3$ over $\mathbb{Q}_{3}$ has non-abelian Galois group isomorphic to $C_{13}\rtimes C_{3}$ ([lmfdb:3.13.12.1]).
Proof of main theorem.
Since abelian case has been shown in Corollary 3.11, we only need to prove it for non-abelian group $G\cong H\rtimes C_{p}$ now. In this case, $C(G)\subseteq H$ and $N_{G}(C_{p})=C(G)\times C_{p}$.
@@
$S(F_{G})=S(\sum_{[G^{\prime}]\subseteq H}\frac{f_{G^{\prime}}}{\#N_{G}(G^{\prime})}+\sum_{[G^{\prime}]:\text{modular}}\frac{f_{G^{\prime}}}{\#N_{G}(G^{\prime})}).$
Note that for $G^{\prime}\subseteq H$, if $N_{G}(G^{\prime})=G$, then the class $[G^{\prime}]$ contains $1$ element; if $N_{G}(G^{\prime})\neq G$, then $N_{G}(G^{\prime})=H$ and the class $[G^{\prime}]$ contains $p$ elements. Either way, if we write the sum over $G^{\prime}\subseteq H$, then the denominator is #$\#G=p\#H$. By Proposition 2.12,
$S(\sum_{[G^{\prime}]\subseteq H}\frac{f_{G^{\prime}}}{\#N_{G}(G^{\prime})})=S(\sum_{G^{\prime}\subseteq H}\frac{f_{G^{\prime}}}{\#G})=\frac{(\#H)^{2}}{\#G}=\frac{\#H}{p}.$
By Lemma 4.1, for modular subgroups $G^{\prime}\subseteq G$, $f_{G^{\prime}}=0$ unless $G^{\prime}\subseteq C(G)\times C_{p}$. Therefore, by Proposition 3.9, assuming existence of crepant resolutions,
$S(\sum_{[G^{\prime}]:\text{modular}}\frac{f_{G^{\prime}}}{\#N_{G}(G^{\prime})})=S(\sum_{C_{p}\subseteq G^{\prime}\subseteq C(G)\times C_{p}}\frac{f_{G^{\prime}}}{p\#C(G)})$
$=\sum_{H^{\prime}\subseteq C(G)}\frac{(p^{2}-1)S(f_{H^{\prime}})}{p\#C(G)}=(p-\frac{1}{p})\#C(G).$
Hence $S(F_{G})=p\#C(G)+\frac{\#H-\#C(G)}{p}=\#\mathrm{Conj}(G)$. ∎
5 Computation of Euler characteristic
The idea of proof of main theorem can also be applied for computation of Euler characteristic in other cases: in [yamamoto2022mass], Yamamoto computed a series of Euler characteristic, when the group has a semidirect product structure of a non-modular abelian normal subgroup and the symmetric group $S_{3}$, in characteristic $3$ and dimension $3$. In this section, we introduce examples where the $p$-Sylow subgroup of $G$ is not cyclic (equivalently, the group $G$ is of infinite representation type in characteristic $p$; it is a result by Higman[higman1954indecomposable]), and see that Euler characteristic can still be computed via wild McKay correspondence. To compute $\mathbf{v}$ function in the examples, we need the theorem below.
Theorem 5.1 ([wood2015mass], Theorem 4.8).
If $G$ acts on $V$ by permutation, then $\mathbf{v}(L)=\frac{1}{2}\mathbf{a}(L)$, where
$\mathbf{a}(L):=\sum_{i=0}^{\infty}\frac{\mathrm{codim}(k^{n})^{G_{i}}}{(G_{0}:G_{i})}$
@@ 5 Computation of Euler characteristic
In the remaining part of this section, $k$ is a finite field of order $q>2$ and characteristic $2$, $A_{4}$ is the alternating group with permutation action on $\mathbb{A}_{k}^{4}$, and $C_{2}^{2}=\{e,(12)(34),(13)(24),(14)(23)\}$ is the normal subgroup consisting of modular elements in $A_{4}$, also acting on $\mathbb{A}^{4}$ by permutations.
@@
For an element $L\in A(C_{2}^{2})$, the corresponding extension is determined by composition of two Artin-Schreier extensions, which are given by two different elements
@@
$\mathbf{v}(L)=\left\{\begin{array}[]{ll}\displaystyle\frac{1}{2}(j+1)+k+1&j>0\\ \displaystyle k+1&j=0\end{array}\right.$
@@
By [wu2010ramification], Theorem 3.11, if $L$ is totally ramified, then the ramification groups of $C_{2}^{2}$ are
@@
$a\in\left\{\begin{array}[]{ll}\displaystyle\bigoplus_{j>0,(2,j)=1}(\gamma k\oplus\gamma^{2}k)t^{-j}&L_{1}/K\text{ is unramified}\\ \displaystyle\bigoplus_{j>0,(6,j)=1}k\gamma^{-j}&L_{1}/K\text{ is ramified}\end{array}\right.$
@@
$\mathbf{v}(L)=\left\{\begin{array}[]{ll}\displaystyle\frac{3}{2}(j+1)&L_{1}/K\text{ is unramified}\\ \displaystyle\frac{1}{2}(j+3)&L_{1}/K\text{ is ramified}\end{array}\right.$
@@
Assume that $L=L_{1}(\alpha,\beta)$, where $L_{1}(\alpha)$ and $L_{1}(\beta)$ are Artin-Schreier extensions corresponding to $a,b\in L_{1}$. Then all the nontrivial intermediate fields of $L/L_{1}$ are $L_{1}(\alpha),L_{1}(\beta)$ and $L_{1}(\alpha+\beta)$. Then any element of order $3$ in $\mathrm{Gal}(L/K)\cong A_{4}$ should give a cyclic permutation on $\alpha,\beta$ and $\alpha+\beta$.
In other words, $L/L_{1}$ is determined by an element $a\in L_{1}$ such that for a generator $\tau\in\mathrm{Gal}(L_{1}/K)$, $a,\tau(a)$ and $\tau^{2}(a)$ are distinct and $\tau^{2}(a)=\tau(a)+a$. Assume that $L_{1}=K(\gamma)$ and $\tau(\gamma)=\zeta_{3}\gamma$, where $\gamma^{3}$ is in $\mathrm{res}(K)$ if $L_{1}/K$ is unramified, while $\gamma$ a uniformizer of $\mathcal{O}_{L_{1}}$ if $L_{1}/K$ is totally ramified. Then computation shows that $a$ can be taken as stated.
Then we compute value of $\mathbf{v}$ function. If $L_{1}/K$ is unramified, then the ramification groups are determined by $C_{2}^{2}$ part as in proof of Proposition 5.2, where $v_{L_{1}}(a)=v_{L_{1}}(\tau(a))=v_{L_{1}}(\tau^{2}(a))=-j$, hence $\mathbf{v}(L)=\frac{3}{2}(j+1)$; if $L_{1}/K$ is ramified, then the ramification groups are
@@
$=6q^{2}(q-1)\frac{1}{1-q^{-1}}+4(q-1)(q-2)q\frac{1}{1-q^{-1}}+12(q-1)^{2}q\frac{1}{1-q^{-1}}\frac{q^{-1}}{1-q^{-1}}$
@@
We use the same way to compute $f_{A_{4}}$. For $L\in A(A_{4})$, it is determined by $L_{1}\in A(C_{3})$ and $a\in L_{1}$ as in Proposition 5.3. Since $L_{1}/K$ is a Kummer extension, we have $1$ choice for unramified one and $3$ choices for totally ramified ones (ignoring $C_{3}$-action on it). By abuse of notation, let $\tau$ be an element in $\mathrm{Gal}(L/K)$ of order $3$, whose restrction on $L_{1}$ generates $\mathrm{Gal}(L_{1}/K)$. For $L/L_{1}$, on one hand, $a,\tau(a),\tau^{2}(a)$ give the same extension; on the other hand, there are $3!=6$ ways to equip the extension with a $C_{2}^{2}$-action. Finally, whenever the $C_{2}^{2}$-extension is determined, one can choose an element of order $3$ in $A_{4}$ from a specific conjugacy class such that it acts on $L$ as $\tau$, which contains $4$ choices.
@@
Although the author does not know if $X_{1}$ has a crepant resolution, $X_{2}$ does have a crepant resolution with Euler characteristic $10$, as constructed in [fan2023crepant].
@@
We can furthermore compute that $F_{C_{2}^{2}}=q^{4}+4q^{3}+q^{2}$ and $F_{A_{4}}=q^{4}+6q^{3}+3q^{2}$, which can be seen as a realization of stringy motives for quotient singularities $X_{1}$ and $X_{2}$ over $\mathbb{F}_{q}$, from the perspective of motivic wild McKay correspondence. This coincides the construction in [fan2023crepant], where the class of crepant resolution of $\mathbb{A}^{4}/A_{4}$ in the (modified) Grothendieck ring is $\mathbb{L}^{4}+6\mathbb{L}^{3}+3\mathbb{L}^{2}$.

### v2

@@ 4 The non-abelian case of the main theorem
Let $G$ be a non-abelian group with a semidirect product structure $H\rtimes C_{p}$ as stated in Proposition 2.19. Then there does not exist any Galois extension $L/K$ such that $\mathrm{Gal}(L/K)\cong G$. In particular, $f_{G}=0$.
@@
If there exists such a Galois extension $L/K$, then $L^{C(G)}/K$ is a Galois extension with its Galois group isomorphic to $G/C(G)\cong(H/C(G))\rtimes C_{p}$. Therefore, we can assume $C(G)=\{e\}$ without loss of generality.
@@
By classical results (one can refer to [10], IV, Corollary 4), $G_{0}$, as a normal subgroup of $G$, is a semidirect product of a normal $p$-subgroup and a non-modular cyclic group. If $G_{0}$ contains an element of order $p$, then $G_{0}$ has to be the whole $G$ since $G_{0}\triangleleft G$, which contradicts the semidirect product structure of $G_{0}$. Hence $C_{p}$ is not contained in $G_{0}$. Since $G_{0}=\{g\in G\mid g\text{ acts trivially on }\mathrm{res}(L)\}$, $C_{p}$ acts effectively on $\mathrm{res}(L)$.
Thus, for $L^{H}/K$, as an Artin-Schreier extension that is either unramified or totally ramified, its residue degree $f_{L^{H}/K}>1$. Consequently, $L^{H}/K$ is unramified. Therefore, if we write the local field $L\cong\mathrm{res}(L)((t_{L}))$, there exists an element $g\in G$ of order $p$ acting trivially on $t_{L}$. We assume that $C_{p}$ is generated by such $g$ without loss of generality.
Take an element $\sigma$ of prime order $l$ ($l\neq p$) from $H$. Then $L/L^{<\sigma>}$ is studied by Kummer theory. If $L/L^{<\sigma>}$ is unramified, then we can similarly assume that $\sigma$ acts trivially on $t_{L}$, and thus $\sigma$ has to commute with $C_{p}$, since the Galois groups of extensions of finite fields are abelian. If $L/L^{<\sigma>}$ is totally ramified, then $\sigma$ acts trivially on $\mathrm{res}(L)$, and hence also commutes with $C_{p}$.
@@
Lemma 4.31 only holds in equal characteristic. In mixed characteristic, for example, the splitting field of $x^{13}+3$ over $\mathbb{Q}_{3}$ has a non-abelian Galois group isomorphic to $C_{13}\rtimes C_{3}$ ([11]).
Proof of the main theorem.
Since the abelian case has been shown in Corollary 3.30, we only need to prove the main theorem for a non-abelian group $G\cong H\rtimes C_{p}$ now. In the non-abelian case, $C(G)\subseteq H$ and $N_{G}(C_{p})=C(G)\times C_{p}$.
@@
$S(F_{G})=S\left(\sum_{[G^{\prime}]:G^{\prime}\subseteq H}\frac{f_{G^{\prime}}}{\#N_{G}(G^{\prime})}+\sum_{[G^{\prime}]:\text{modular}}\frac{f_{G^{\prime}}}{\#N_{G}(G^{\prime})}\right).$
Note that for $G^{\prime}\subseteq H$, if $N_{G}(G^{\prime})=G$, then the class $G^{\prime}]$ contains one element; if $N_{G}(G^{\prime})\neq G$, then $N_{G}(G^{\prime})=H$ and the class $G^{\prime}]$ contains $p$ elements. Either way, if we write the sum over $G^{\prime}\subseteq H$, then the denominator is #G=p#H\#G=p\#H. By Proposition 2.17,
$S\left(\sum_{[G^{\prime}]:G^{\prime}\subseteq H}\frac{f_{G^{\prime}}}{\#N_{G}(G^{\prime})}\right)=S\left(\sum_{G^{\prime}\subseteq H}\frac{f_{G^{\prime}}}{\#G}\right)=\frac{(\#H)^{2}}{\#G}=\frac{\#H}{p}.$
By Lemma 4.31, for modular subgroups $G^{\prime}\subseteq G$, $f_{G^{\prime}}=0$ unless $G^{\prime}\subseteq C(G)\times C_{p}$. Therefore, by Proposition 3.28, assuming the existence of crepant resolutions,
$S\left(\sum_{[G^{\prime}]:\text{modular}}\frac{f_{G^{\prime}}}{\#N_{G}(G^{\prime})}\right)=S\left(\sum_{C_{p}\subseteq G^{\prime}\subseteq C(G)\times C_{p}}\frac{f_{G^{\prime}}}{p\#C(G)}\right)$
$=\sum_{H^{\prime}\subseteq C(G)}\frac{(p^{2}-1)S(f_{H^{\prime}})}{p\#C(G)}=\left(p-\frac{1}{p}\right)\#C(G).$
Hence
$S(F_{G})=p\#C(G)+\frac{\#H-\#C(G)}{p}=\#\mathrm{Conj}(G).$
5 Computation of the Euler characteristic
The idea of the proof of our main theorem can also be applied to compute the Euler characteristic in other cases: in [16], Yamamoto computed a series of Euler characteristics, when the group has a semidirect product structure of a non-modular abelian normal subgroup and the symmetric group $S_{3}$, in characteristic $3$ and dimension $3$. In this section, we introduce two examples where the $p$-Sylow subgroup of $G$ is not cyclic (equivalently, the group $G$ is of infinite representation type in characteristic $p$; this is a result by Higman[7]), and see that the Euler characteristic can still be computed via the wild McKay correspondence. To compute the $\mathbf{v}$ function in the examples, we need the theorem below.
Theorem 5.33 ([13], Theorem 4.8).
Let $G$ be a finite group acting on a representation space $V$. The fields $k,K,L$ and the $\mathbf{v}$ function are as defined in Section 2. If $G$ acts on $V$ by permutation, then $\mathbf{v}(L)=\frac{1}{2}\mathbf{a}(L)$, where
$\mathbf{a}(L):=\sum_{i=0}^{\infty}\frac{\mathrm{codim}(k^{n})^{G_{i}}}{[G_{0}:G_{i}]}$
@@ 5 Computation of the Euler characteristic
In the remaining part of this section, $k$ is a finite field of order $q>2$ and characteristic $2$, $A_{4}$ is the alternating group with the permutation action on $\mathbb{A}_{k}^{4}$, and $C_{2}^{2}=\{e,(12)(34),(13)(24),(14)(23)\}$ is the normal subgroup containing modular elements in $A_{4}$, also acting on $\mathbb{A}^{4}$ by permutation.
@@
For an element $L\in A(C_{2}^{2})$, the corresponding extension is determined by the composition of two Artin-Schreier extensions, which are given by two different elements
@@
$\mathbf{v}(L)=\left\{\begin{array}[]{ll}\displaystyle\frac{1}{2}(j+1)+k+1&j>0\\ \displaystyle k+1&j=0\end{array}\right.$.
@@
By [14], Theorem 3.11, if $L$ is totally ramified, then the ramification groups of $C_{2}^{2}$ are
@@
$a\in\left\{\begin{array}[]{ll}\displaystyle\bigoplus_{j>0,(2,j)=1}(\gamma k\oplus\gamma^{2}k)t^{-j}&L_{1}/K\text{ is unramified}\\ \displaystyle\bigoplus_{j>0,(6,j)=1}k\gamma^{-j}&L_{1}/K\text{ is ramified}\end{array}\right.$.
@@
$\mathbf{v}(L)=\left\{\begin{array}[]{ll}\displaystyle\frac{3}{2}(j+1)&L_{1}/K\text{ is unramified}\\ \displaystyle\frac{1}{2}(j+3)&L_{1}/K\text{ is ramified}\end{array}\right.$.
@@
Assume that $L=L_{1}(\alpha,\beta)$, where $L_{1}(\alpha)$ and $L_{1}(\beta)$ are Artin-Schreier extensions corresponding to $a,b\in L_{1}$. All the nontrivial intermediate fields of $L/L_{1}$ are $L_{1}(\alpha),L_{1}(\beta)$ and $L_{1}(\alpha+\beta)$. Then any element of order $3$ in $\mathrm{Gal}(L/K)\cong A_{4}$ should give a cyclic permutation on $\alpha,\beta$ and $\alpha+\beta$.
In other words, $L/L_{1}$ is determined by an element $a\in L_{1}$ such that for a generator $\tau\in\mathrm{Gal}(L_{1}/K)$, $a,\tau(a)$ and $\tau^{2}(a)$ are distinct and $\tau^{2}(a)=\tau(a)+a$. Assume that $L_{1}=K(\gamma)$ and $\tau(\gamma)=\zeta_{3}\gamma$, where $\gamma^{3}$ is in $\mathrm{res}(K)$ if $L_{1}/K$ is unramified, or $\gamma$ is a uniformizer of $\mathcal{O}_{L_{1}}$ if $L_{1}/K$ is totally ramified. Then an easy computation shows that $a$ can be taken as stated in the proposition.
Then we compute the value of the $\mathbf{v}$ function. If $L_{1}/K$ is unramified, then the ramification groups are determined by the $C_{2}^{2}$ part as in the proof of Proposition 5.34, where $v_{L_{1}}(a)=v_{L_{1}}(\tau(a))=v_{L_{1}}(\tau^{2}(a))=-j$, hence $\mathbf{v}(L)=\frac{3}{2}(j+1)$; if $L_{1}/K$ is ramified, then the ramification groups are
@@
$=6q^{2}(q-1)\frac{1}{1-q^{-1}}+4(q-1)(q-2)q\frac{1}{1-q^{-1}}$
$+12(q-1)^{2}q\frac{1}{1-q^{-1}}\frac{q^{-1}}{1-q^{-1}}$
@@
We compute $f_{A_{4}}$ in the same way. For $L\in A(A_{4})$, it is determined by $L_{1}\in A(C_{3})$ and $a\in L_{1}$ as in Proposition 5.35. Since $L_{1}/K$ is a Kummer extension, we have $1$ choice for the unramified one and $3$ choices for the totally ramified ones (ignoring the $C_{3}$-action on it). By abuse of notation, let $\tau$ be an element of $\mathrm{Gal}(L/K)$ of order $3$, whose restriction on $L_{1}$ generates $\mathrm{Gal}(L_{1}/K)$. For $L/L_{1}$, on one hand, $a,\tau(a),\tau^{2}(a)$ give the same extension; on the other hand, there are $3!=6$ ways to equip the extension with a $C_{2}^{2}$-action. Finally, whenever the $C_{2}^{2}$-extension is determined, one can choose an element of order $3$ in $A_{4}$ from a specific conjugacy class such that it acts on $L$ as $\tau$, which contains $4$ choices.
@@
Although the author does not know if $X_{1}$ has a crepant resolution, $X_{2}$ does have a crepant resolution with Euler characteristic $10$, as constructed in [5].
@@
We can furthermore compute that $F_{C_{2}^{2}}=q^{4}+4q^{3}+q^{2}$ and $F_{A_{4}}=q^{4}+6q^{3}+3q^{2}$, which can be seen as a realization of the stringy motives for quotient singularities $X_{1}$ and $X_{2}$ over $\mathbb{F}_{q}$, from the perspective of the motivic wild McKay correspondence. This coincides with the construction in [5], where the class of the crepant resolution of $\mathbb{A}^{4}/A_{4}$ in the (modified) Grothendieck ring is $\mathbb{L}^{4}+6\mathbb{L}^{3}+3\mathbb{L}^{2}$.
Acknowledgements
The author expresses his gratitude to Prof. Takehiko Yasuda, Prof. Dr. Christian Liedtke, and Dr. Rin Gotou for valuable discussions. The author thanks his supervisor, Prof. Yukari Ito, for her advice and consideration. The author also thanks the anonymous referees for their helpful comments. This work was supported by JSPS KAKENHI Grant Number JP25KJ1153.